Solve the following inequality: -2x^(2/3)>(or equal to)6
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

Solve the following inequality: -2x^(2/3)>(or equal to)6
---
-2x^(2/3)>= 6
---
x^(2/3) >= -3
---
(x^(2/3))^3 = (-3)^3
x^2 = -27
---
x = 3sqrt(3)i
=================
Cheers,
Stan H.
=================

Answer by lwsshak3(11628)

Solve the following inequality: -2x^(2/3)>(or equal to)6
**
-2x^(2/3)≥6
divide by -2
x^2/3≥6/-2≥-3
raise both sides to 3/2 power
x≥3^3/2=√27
solution: [√27,∞)